# Beatty Sequences for a Quadratic Irrational: Decidability and Applications

### 24 Mar 2024 v2

@@ 6 Beatty sequences in Walnut
We now explain how to use Beatty sequences of quadratic irrationals in Walnut. Suppose $\alpha$ and $\beta$ lie in a quadratic field $\mathbb{Q}(\gamma)$, where $\gamma$ is a purely periodic continued fraction.
If the period of $\gamma$’s continued fraction starts with $1$, rotate the period to get $\gamma^{\prime}$ with a period that begins with a number larger than $1$. Then $\gamma^{\prime}$ generates the same field as $\gamma$. This is necessary because Walnut assumes that the continued fraction represents a number in the open interval $(0,{1\over 2})$. Rotation of the period only fails when $\gamma=[0,1,1,1,\ldots]$, but in that case we can just use the Fibonacci numeration system instead, based on the real number $[0,2,1,1,1,\ldots]$.
@@
Next, we obtain the synchronized automaton for $(\lfloor n\gamma\rfloor)_{n\geq 1}$. In Theorem 4, we have $q_{m}=4$ and $q_{m-1}=3$. We use the following Walnut commands:
@@
We should say something briefly about the syntax of Walnut here. The symbol & represents logical AND, ~ represents logical NOT, | represents logical OR, => is implication, <=> is IFF, E is the existential quantifier, and A is the universal quantifier. The ost command creates an Ostrowski numeration system, the def command defines an automaton for later use, the reg command allows defining a regular expression, and the eval command evaluates a formula with no free variables and reports TRUE or FALSE. We can multiply and do integer division by natural number constants.
@@ 7 The additive theory of Beatty sequences
By a classical theorem (see Nathanson [39, p. 366, Theorems 11.6 and 11.7]) every Beatty sequence forms an asymptotic additive basis where an upper bound on the number of terms needed to represent every sufficiently large number is (effectively) computable. Now we just write first-order statements that every sufficiently large integer is a sum of $h$ terms of the Beatty sequence, and evaluate them using the algorithm of Theorem 2, until we find the smallest $h$. The algorithm is guaranteed to terminate by Nathanson’s theorem.
Similarly we can write a first-order statement that every natural number is a sum of some number of terms. ∎
@@
Let $\varphi=(1+\sqrt{5})/2$, the golden ratio, and consider the sequence ${\bf c}=(n\varphi+{1\over 2})_{n\geq 0}$. This is sequence A007067 in the OEIS [46], and was originally studied by Johann Bernoulli [47, pp. 66–69].
@@
The sequence $\bf c$ is $\varphi$-synchronized.
@@
Starting from the known automaton (called phin) for the Beatty sequence $(\lfloor n\varphi\rfloor)_{n\geq 0}$ (see, e.g., [45, §10.11]), we can get an automaton for $\bf c$ as follows (in Walnut)
@@
This produces the synchronized automaton for $\bf c$ in Figure 2.
@@
Every integer $n\not=1$ can be written as the sum of two elements of $\bf c$.
@@ 8 Proofs of some conjectures of Reble
We implement the formulas above in Walnut. Here F is Walnut’s abbreviation for the Fibonacci word, indexed starting at position 00.
@@ 9 A sequence of Hildebrand, Li, Li, and Xie
combine diff has11=-1 diff1=1 diff2=2:
which gives us the automaton in Figure 3, once the dead state corresponding to invalid Fibonacci representations is removed. ∎
@@ 10 The Kimberling swappage problem
Let $L(n)=\lfloor n\varphi\rfloor$ be the lower Wythoff sequence and $U(n)=\lfloor n\varphi^{2}\rfloor$ be the upper Wythoff sequence. Kimberling defined the “lower even swappage sequence” $V_{\rm le}(n)$ as follows: if $U(n)$ is even, then $V_{\rm le}(n):=U(n)$. Otherwise, if $U(n)$ is odd, then $V_{\rm le}(n):=L(i)$, where $i$ is the least index such that $U(n-1)<L(i)<U(n+1)$ and $L(i-1)<U(n)<L(i+1)$. Here $(V_{\rm le}(n))$ is OEIS sequence A141104. He conjectured that the sequence $S(n):=V_{\rm le}(n)/2$ consists of those positive integers not in the Beatty sequence $(\lfloor n\varphi^{3}\rfloor)_{n\geq 1}$, which was later proved by Russo and Schwiebert [42].
We can use Walnut to provide an alternate proof of Kimberling’s conjecture. The ideas are those similar to those of the previous section. First, using the ideas in [45, §10.15], we guess a synchronized automaton leswap for $V_{\rm le}(n)$ (it has $34$ states), and we then verify it satisfies the definition.
@@
eval kimber "?msd_fib Ax (x>0) => (En (n>=1) & \$phi3n(n,x))
   <=> (~En,y (n>=1) & \$leswap(n,y) & 2*x=y)":
@@ 11 Sums-complement sets
Recall that a set $S$ is said to be $\gamma$-Ostrowski automatic if there is a finite automaton accepting the $\gamma$-Ostrowski representation of $n$ as input and accepting if and only if $n\in S$. Then we have the following proposition.
@@
Suppose $10<\gamma<$ 1 is a quadratic irrational with purely periodic continued fraction
$[0,a_{1},a_{2},\ldots,a_{m},a_{1},a_{2},\ldots,a_{m},a_{1},a_{2},\ldots,a_{m},\ldots],$
@@ 12 Sequences related to $\sqrt{2}$
We start with A097508$(n)$, whose definition is $\lfloor n(\sqrt{2}-1)\rfloor$. In the corresponding Ostrowski numeration system, the digits can be 0, 1 or 2. The shifting automaton shift1 can be built using Corollary 5, and it is depicted in Figure 4.
@@
Next we turn to OEIS sequence A001951$(n)$, defined by $\lfloor n\sqrt{2}\rfloor$. We apply Theorem 7 using $a=1$, $b=1$, $c=1$, $d=0$, and $e=0$ to create the automaton for A001951 using the following Walnut command.
@@
Next we turn to OEIS sequence A003151$(n)$, defined by $\lfloor n(\sqrt{2}+1)\rfloor$. We apply Theorem 7 using $a=2$, $b=1$, $c=1$, $d=0$, and $e=0$ to create the automaton for A003151 using the following Walnut command.
@@
Next we look at OEIS sequence A276862$(n)$, defined to be the first difference of A003151. We can create a synchronized automaton for it as follows:
@@
Next we turn to OEIS sequence A080754$(n)$, which is defined to be $\lceil n(\sqrt{2}+1)\rceil$.
@@
$a(t)=\begin{cases}3,&\text{if $t=1$};\\ 2,&\text{if $t=0,2$};\\ a(n),&\text{if $t=a(1)+\cdots+a(n)$};\\ 2,&\text{if $a(t-1)=3$ and $\neg\exists n\geq 0\ \ t=a(1)+\cdots+a(n)$};\\ 3,&\text{if $a(t-1)=2$ and $\neg\exists n\geq 0\ \ t=a(1)+\cdots+a(n)$}.\end{cases}$ (5)
@@
eval check2 "?msd_s2 At (\$b(t-1,3) & ~En \$a080754(n+2,t+5)) => \$b(t,2)":
eval check3 "?msd_s2 At (\$b(t-1,2) & ~En \$a080754(n+2,t+5)) => \$b(t,3)":

### 02 Apr 2026 v3

@@ 6 Beatty sequences in Walnut
[…]
The ost command creates an Ostrowski numeration system, the def command defines an automaton for later use, the reg command allows defining a regular expression, and the eval command evaluates a formula with no free variables and reports TRUE or FALSE. In Walnut, we can multiply and do integer division by natural number constants, but not variables. When an automaton is defined using def, the user supplies a name for it. The automaton can then be used later by invoking its name prefixed with a dollar sign.
If a first-order statement has no free variables, then Walnut returns either TRUE or FALSE, and its computations provide a proof of the result. Otherwise it returns an automaton accepting the values of the free variables that make the logical statement evaluate to TRUE.
The combine command allows the user to combine different automata into a single DFAO with specified outputs.
We now explain how to use Beatty sequences of quadratic irrationals in Walnut. Suppose $\alpha$ and $\beta$ lie in a quadratic field $\mathbb{Q}(\gamma)$, where $1/\gamma$ has a purely periodic continued fraction.
If the period of $\gamma$’s continued fraction starts with $1$, rotate the period to get $\gamma^{\prime}$ with a period that begins with a number larger than $1$. Then $\gamma^{\prime}$ generates the same field as $\gamma$. This is necessary because Walnut assumes that the continued fraction represents a number in the open interval $(0,{1\over 2})$. Rotation of the period only fails when $\gamma=[0,1,1,1,\ldots]$, but in that case we can just use the Fibonacci numeration system instead, based on the real number $0,2,1,1,1,\ldots]$.
@@
States correspond to two consecutive Ostrowski digits that are permitted by the inequalities given in Section 3.
Next, we obtain the synchronized automaton for $(\lfloor n\gamma\rfloor)_{n\geq 1}$. In Theorem 5, we have $q_{P}=4$ and $q_{P-1}=3$. We use the following Walnut commands:
@@
There are obviously some limits to our approach. For example, as the length of the repeating portion of the continued fraction gets longer, the size of the automata grow rapidly, so that working with periods of length (say) longer than $10$ might require computational resources beyond one’s computer.
Another limitation involves theorems about Beatty sequences concerning two numbers that are not in the same quadratic field. There is no obvious way to do this using Walnut.
@@ 7 The additive theory of Beatty sequences
The first claim follows immediately from a classical theorem (see Nathanson [41, p. 366, Theorems 11.6 and 11.7]). Furthermore, the proof of this classical result given in the cited reference shows that every Beatty sequence forms an asymptotic additive basis where an upper bound on the number of terms needed to represent every sufficiently large number is (effectively) computable. Now we just write first-order statements that every sufficiently large integer is a sum of $h$ terms of the Beatty sequence, and evaluate them using the algorithm of Theorem 2, until we find the smallest $h$ that works. The algorithm is then guaranteed to terminate by Nathanson’s theorem.
Similarly we can write a first-order statement that every natural number (not just every sufficiently large natural number) is a sum of some number of terms. ∎
@@
Let $\varphi=(1+\sqrt{5})/2$, the golden ratio, and consider the sequence ${\bf c}=(\lfloor n\varphi+{1\over 2}\rfloor)_{n\geq 0}$. This is sequence A007067 in the OEIS [49], and was originally studied by Johann Bernoulli [50, pp. 66–69]. For convenience, we used the Fibonacci numeration system, which is a feature of Walnut.
@@
The sequence $\bf c$ is Fibonacci synchronized.
@@
Starting from the known automaton (called phin) for the Beatty sequence $(\lfloor n\varphi\rfloor)_{n\geq 0}$ (see, e.g., [48, §10.11]), we can get an automaton for $\bf c$ as follows (in Walnut)
@@
This produces the synchronized automaton for $\bf c$ in Figure 2. The reg command creates an automaton for a regular expression that implements the assertion that the second argument is the left-shift of the first argument.
@@
Every non-negative integer $n\not=1$ can be written as the sum of two elements of $\bf c$.
@@
Using the same ideas we can prove the following. Call two sets of integers $S,T$ asymptotically additively equivalent if there exists an integer $M$ such that every integer $n\geq M$ can be written as the sum of two elements of $S$ if and only if the same is true of $T$.
Theorem 16.
Given $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}$ real numbers in the same quadratic field, we can decide if the two inhomogeneous Beatty sequences $(\lfloor\alpha_{1}n+\beta_{1}\rfloor)_{n\geq 1}$ and $(\lfloor\alpha_{2}n+\beta_{2}\rfloor)_{n\geq 1}$ are asymptotically additively equivalent.
The (simple) proof is left to the reader.
@@ 8 Proofs of some conjectures of Reble
We implement the formulas above in Walnut. Here F is Walnut’s abbreviation for the Fibonacci word, indexed starting at position 0, and phin is the synchronized automaton computing $\lfloor\varphi n\rfloor$ for $\varphi=(1+\sqrt{5})/2$, the golden ratio, given in the proof of Theorem 14.
@@
9 Graham’s question
For real numbers $\alpha>0$ and $\beta$, define the set $S(\alpha,\beta)=\{\lfloor\alpha n+\beta\rfloor\,:\,n\geq 1\}$. Graham [22] studied the following question: for which finite sets of pairs of real numbers $\{(\alpha_{i},\beta_{i})\,:\,1\leq i\leq r\}$ does every sufficiently large integer belong to exactly one $S(\alpha_{i},\beta_{i})$? Such a finite set of pairs is called an eventual covering family or ECF.
Our results prove that a special case is decidable:
Theorem 18.
Suppose that all the $\alpha_{i}$ and $\beta_{i}$, $1\leq i\leq r$, lie in the same quadratic field. Then Graham’s problem is decidable for $\{(\alpha_{i},\beta_{i})\,:\,1\leq i\leq r\}$.
Proof.
We can create a first-order logical formula asserting that there exists an $M$ such that, for all $n\geq M$, we have
  • (i)
    there exist $k$ and at least one $i$, $1\leq i\leq r$, such that $n=\lfloor\alpha_{i}k+\beta_{i}\rfloor$;
  • (ii)
    there do not exist $i,j,k_{i},k_{j}$, $1\leq i<j\leq r$, such that $n=\lfloor\alpha_{i}k_{i}+\beta_{i}\rfloor=\lfloor\alpha_{j}k_{j}+\beta_{j}\rfloor$.
It now follows from our approach that both of these are decidable. ∎
Example 19.
We show that $\{(2\alpha,0),(2\alpha,\alpha),(\alpha^{2},0)\}$ is such a family, where $\alpha=(1+\sqrt{5})/2$.
eval case_i "?msd_fib An (n>=2) => Ek k>=1 & (\$phin(2*k,n) | \$phin(2*k+1,n)
   | \$phi2n(k,n))":
eval case_ii "?msd_fib ~En,j,k j>=1 & k>=1 & ((\$phin(2*j,n) & \$phin(2*k+1,n))
   | (\$phin(2*j+1,n) & \$phi2n(k,n)) | (\$phi2n(j,n) & \$phin(2*k,n)))":
Both of these return TRUE.
@@ 10 A sequence of Hildebrand, Li, Li, and Xie
combine diff diff1=1 diff2=2:
which gives us the automaton in Figure 3. ∎
@@ 11 The Kimberling swappage problem
Let $L(n)=\lfloor n\varphi\rfloor$ be the lower Wythoff sequence and $U(n)=\lfloor n\varphi^{2}\rfloor$ be the upper Wythoff sequence. Kimberling defined the “lower even swappage sequence” $V_{\rm le}(n)$ as follows: if $U(n)$ is even, then $V_{\rm le}(n):=U(n)$. Otherwise, if $U(n)$ is odd, then $V_{\rm le}(n):=L(i)$, where $i$ is the least index such that U(n1)<L(i)<U(n+1)U(n-1)<L(i)<U(n+1) and L(i1)<U(n)<L(i+1)L(i-1)<U(n)<L(i+1). Here $(V_{\rm le}(n))$ is OEIS sequence A141104. He conjectured that the sequence $S(n):=V_{\rm le}(n)/2$ consists of those positive integers not in the Beatty sequence $(\lfloor n\varphi^{3}\rfloor)_{n\geq 1}$, which was later proved by Russo and Schwiebert [45].
We can use Walnut to provide an alternate proof of Kimberling’s conjecture. The ideas are those similar to those of the previous section. First, using the ideas in [48, §10.15], we guess a synchronized automaton leswap for $V_{\rm le}(n)$ (it has $34$ states), and we then verify it satisfies the definition. Here phi2n is a synchronized automaton computing $\lfloor\varphi^{2}n\rfloor$, taken from [48, §10.11].
@@
The two results of TRUE from the last two commands verify that our guessed sequence satisfies the definition.
@@
def phi3n "?msd_fib Ey \$phin(2*n,y) & z=y+n":
eval kimber "?msd_fib Ax (x>0) => ((En (n>=1) & \$phi3n(n,x))
   <=> (~En,y (n>=1) & \$leswap(n,y) & 2*x=y))":
@@ 12 Sums-complement sets
A set $S$ is said to be $\gamma$-Ostrowski automatic if there is a finite automaton accepting the $\gamma$-Ostrowski representation of $n$ as input and accepting if and only if $n\in S$. Then we have the following proposition.
@@
Suppose $0<\gamma<1$ is a quadratic irrational such that $1\over\gamma$ has purely periodic continued fraction
$a_{1},a_{2},\ldots,a_{P},a_{1},a_{2},\ldots,a_{P},a_{1},a_{2},\ldots,a_{P},\ldots],$
@@ 13 Sequences related to $\sqrt{2}$
We start with A097508(n)(n), whose definition is $\lfloor n(\sqrt{2}-1)\rfloor$. In the corresponding Ostrowski numeration system, the digits can be 0, 1 or 2. The shifting automaton shift1 can be built using Corollary 6, and it is depicted in Figure 4.
@@
Next we turn to OEIS sequence A001951(n)(n), defined by $\lfloor n\sqrt{2}\rfloor$. We apply Theorem 9 using $a=1$, $b=1$, $c=1$, $d=0$, and $e=0$ to create the automaton for A001951 using the following Walnut command.
@@
Next we turn to OEIS sequence A003151(n)(n), defined by $\lfloor n(\sqrt{2}+1)\rfloor$. We apply Theorem 9 using $a=2$, $b=1$, $c=1$, $d=0$, and $e=0$ to create the automaton for A003151 using the following Walnut command.
@@
Next we look at OEIS sequence A276862(n)(n), defined to be the first difference of A003151. We can create a synchronized automaton for it as follows:
@@
And Walnut returns TRUE.
Next we turn to OEIS sequence A080754(n)(n), which is defined to be $\lceil n(\sqrt{2}+1)\rceil$.
@@
[…]
@@
eval check2 "?msd_s2 At (\$b2(t-1,3) & ~En \$a080754(n+2,t+5)) => \$b2(t,2)":
eval check3 "?msd_s2 At (\$b2(t-1,2) & ~En \$a080754(n+2,t+5)) => \$b2(t,3)":
@@
14 Fractional parts
Let $\{x\}$ denote the fractional part of $x$, that is, ${x\bmod 1}=x-\lfloor x\rfloor$. Inspired by recent work of Khani, Valizadeh, and Zarei [32], we show how to handle assertions like $\{\alpha x+\beta\}<\{\alpha^{\prime}y+\beta^{\prime}\}$ in Walnut, where $x$ and $y$ are natural number variables and $\alpha,\alpha^{\prime},\beta,\beta^{\prime}$ all lie in the same quadratic field. The key idea is the next very simple proposition.
Proposition 24.
Let $x,y$ be real numbers. Then $\{x\}<\{y\}$ if and only if $\lfloor x-y\rfloor<\lfloor x\rfloor-\lfloor y\rfloor$.
Proof.
Write $x=\lfloor x\rfloor+\{x\}$, $y=\lfloor y\rfloor+\{y\}$. Then
$\{x\}<\{y\}$ $\iff x-\lfloor x\rfloor<y-\lfloor y\rfloor$
$\iff x-y<\lfloor x\rfloor-\lfloor y\rfloor$
$\iff\lfloor x-y\rfloor<\lfloor x\rfloor-\lfloor y\rfloor.$
The virtue of Proposition 24 is that we can turn a first-order logic statement involving comparison of the sizes of fractional parts into a first-order statement involving floors.
Example 25.
Let us create an automaton to decide if $\{\alpha x\}<\{\alpha y\}$, where $\alpha=(1+\sqrt{5})/2$, the golden ratio. Once more we use the automaton phin that accepts the Zeckendorf representation of the natural number pair $(x,y)$ in parallel if and only if $y=\lfloor\alpha x\rfloor$.
Translation of Proposition 24 into Walnut involves a tiny bit of extra effort, because the default domain for integers is $\mathbb{N}$ and not $\mathbb{Z}$. Thus we need two cases, depending on whether $x>y$ or $x<y$.
reg shift {0,1} {0,1} "([0,0]|[0,1][1,1]*[1,0])*":
def phin "?msd_fib (x=0 & y=0) | Ez \$shift(x-1,z) & y=z+1":
def cfp "?msd_fib Er,s,t \$phin(x,r) & \$phin(y,s) & \$phin(x-y,t) & t+s<r":
# decides if {alpha*x} < {alpha*y} provided x>y
def compare "?msd_fib (x>y & \$cfp(x,y)) | (y>x & ~\$cfp(y,x))":
This generates an automaton compare of 22 states. The inputs are $x$ and $y$ in Zeckendorf representation, most-significant-digit first, and the input is accepted if and only if $\{\alpha x\}<\{\alpha y\}$.
This implementation works for $x,y\geq 0$. But we might want to also be able to compare $\{\alpha x\}$ and $\{\alpha y\}$ where one or both of $x,y$ is negative. We can do this using Proposition 24 together with the identity $\lfloor-x\rfloor=\lceil x\rceil$, and the fact that $\lceil x\rceil=\lfloor x\rfloor+1$ for $x\not\in\mathbb{Z}$.
Example 26.
Similarly we can check if $\{\alpha x\}<c/d$ for quadratic $\alpha$ and fixed natural number constants $0\leq c<d$. By Proposition 24 and Lemma 4 we have
$\{\alpha x\}<c/d\iff\left\lfloor{{{\lfloor\alpha dx\rfloor}-c}\over{d}}\right\rfloor<\lfloor\alpha x\rfloor.$
For example, for $\alpha=(1+\sqrt{5})/2$ and $c=1$, $d=2$, we get an automaton to check if $\{\alpha x\}<1/2$, as follows:
def cmphalf "?msd_fib (x=0) | Ey,z \$phin(2*x,y) & \$phin(x,z) & (y-1)/2 < z":
This gives an $8$-state automaton to decide if $\{\alpha x\}<1/2$.
Dan Asimov published a query on the math-fun mailing list on September 23 2025, which we summarize as follows: suppose $\alpha$ is an irrational number, and $x,y,z$ are three integers. Mapping the half-open interval $[0,1)$ to the unit circle in the standard way via $z\rightarrow\exp(2\pi iz)$, how do we determine the relative order of the fractional parts $\{\alpha x\}$, $\{\alpha y\}$, $\{\alpha z\}$ on the circle, clockwise or counter-clockwise? He was particularly interested in the case $\alpha=\sqrt{2}$.
We can do this for any quadratic irrational using the ideas of this section, because the relative order depends completely on the three pairwise comparisons of the fractional parts of $\{\alpha x\}$, $\{\alpha y\}$, and $\{\alpha z\}$.
We now illustrate this idea for $\sqrt{2}$. From Section 13 we have an automaton a001951 that computes the map $n\rightarrow\lfloor\sqrt{2}n\rfloor$.
def cfp2 "?msd_s2 Er,s,t \$a001951(x,r) & \$a001951(y,s) & \$a001951(x-y,t) & t+s<r":
def compare2 "?msd_s2 (x>y & \$cfp2(x,y)) | (y>x & ~\$cfp2(y,x))":
def order2 "?msd_s2 \$compare2(x,y) & \$compare2(y,z)":
def ccw "?msd_s2 \$order2(x,y,z) | \$order2(y,z,x) | \$order2(z,x,y)":
This gives us a $44$-state automaton ccw that solves the problem, accepting if the ordering for input $(x,y,z)$ is counter-clockwise and rejecting otherwise.
15 The converse to Theorem 9
In this section we prove the converse to Theorem 9.
We need a lemma. Recall that the subword complexity (aka factor complexity) of a word is the function $\rho(n)$ counting the number of distinct length-$n$ factors.
Lemma 27.
Suppose $0<\alpha_{i},\beta_{i}<1$ are real numbers for $1\leq i\leq k$. Let
$1,\alpha_{1},\ldots,\alpha_{k}$
Proof.
By the multidimensional version of Kronecker’s theorem (see, e.g., [24, Thm. 442]), the linear independence criterion implies that for $\epsilon>0$ and all choices of the $b_{i}$, there exist integers $q$ and $p_{i}$, $1\leq i\leq k$ such that $|\sum_{1\leq i\leq k}q\alpha_{i}-p_{i}+b_{i}|<\epsilon$.
We can now prove the converse to Theorem 9.
Theorem 28.
Suppose $0<\gamma<1$ is a quadratic irrational such that $1/\gamma$ has a purely periodic continued fraction expansion. Suppose the Beatty sequence $(\lfloor n\alpha+\beta\rfloor)_{n\geq 1}$ is $\gamma$-Ostrowski synchronized. Then $\alpha$ and $\beta$ both belong to $\mathbb{Q}(\gamma)$. Furthermore, the expression of both $\alpha$ and $\beta$ in terms of $\gamma$ is computable.
Proof.
Suppose $(\lfloor n\alpha+\beta\rfloor)_{n\geq 1}$ is $\gamma$-Ostrowski synchronized. Our first goal is to prove that $\alpha\in\mathbb{Q}(\gamma)$. Suppose $1,\alpha,\gamma$ are linearly independent over $\mathbb{Q}$. If the Beatty sequence $(\lfloor n\alpha+\beta\rfloor)_{n\geq 1}$ is $\gamma$-Ostrowski synchronized, then the corresponding Sturmian sequence (possibly recoded over the alphabet $\{0,1\}$ if necessary) is $\gamma$-Ostrowski automatic. This is also true for the Sturmian sequence associated with slope $\gamma$ and intercept 0. Then the product of these two sequences (over the alphabet $\{0,1\}^{2}$) is also $\gamma$-Ostrowski automatic. Hence this product has subword complexity $\Theta(n^{2})$, but from Corollary 3, we know that the subword complexity of every automatic sequence is $O(n)$.
This contradiction shows that $1,\alpha,\gamma$ are linearly dependent over $\mathbb{Q}$, and hence $\alpha\in\mathbb{Q}(\gamma)$. Therefore there are integers $b,c,d$ with $d\geq 1$ such that that $\alpha=(b\gamma+c)/d$.
We now show that we can computably determine $b,c$, and $d$. To do so we iterate over all possible integer triples by increasing order of $|b|+|c|+d$. For each triple $(b,c,d)$ we construct an $\gamma$-Ostrowski automaton for the Sturmian word with slope $(b\gamma+c)/d$ and intercept 0. We then check (using the logical approach) whether the set of factors occurring in Sturmian word is the same as that for the Sturmian word with slope $\alpha$ and intercept $\beta$. Since all Sturmian words with the same slope have the same set of factors, we eventually find a suitable triple. Then $\alpha=(b\gamma+c)/d$ with $b,c,d$ known integers.
We now turn to discussing $\beta$. If $\beta=0$ there is nothing more to prove. Otherwise, without loss of generality, we may assume that $0<\beta<1$. Put $x=n\alpha$ and $y=-\beta$ in Proposition 24. Then we get that $\{n\alpha\}<\{-\beta\}$ if and only if
$\lfloor n\alpha+\beta\rfloor$ $=\lfloor n\alpha\rfloor$
$=\left\lfloor{{bn\gamma+cn}\over d}\right\rfloor$ (7)
$=\left\lfloor{{\lfloor bn\gamma\rfloor+cn}\over d}\right\rfloor.$
By our hypothesis there is a computable $\gamma$-Ostrowski automaton $A_{1}$ taking $n\geq 1$ as input and computing $\lfloor n\alpha+\beta\rfloor$. Similarly, since we know $b,c,d$ we can construct a $\gamma$-Ostrowski automaton taking $n\geq 1$ as input and computing
$\left\lfloor{{\lfloor bn\gamma\rfloor+cn}\over d}\right\rfloor.$
It follows that there is an automaton $A_{2}$ that takes $n$ as input and checks if $\{n\alpha\}<\{-\beta\}$.
Now make a new automaton $A_{3}$ that accepts those integers $m\geq 1$ such that $m\equiv 0\ \mbox{\rm(mod}\ d\mbox{\rm)}$ and
$m\equiv 0\ \mbox{\rm(mod}\ d\mbox{\rm)}\quad\text{ and }\quad\{m\alpha\}<\{-\beta\}.$ (8)
and further there is no $m^{\prime}$, $0<m^{\prime}<m$, satisfying Eq. (8) such that $\{m\alpha\}<\{m^{\prime}\alpha\}$. Roughly speaking, $A_{3}$ accepts those $m\equiv 0\ \mbox{\rm(mod}\ d\mbox{\rm)}$ such that $\{m\alpha\}$ consists of better and better approximations to $\{-\beta\}$. Again by Kronecker’s theorem the automaton $A_{3}$ accepts infinitely many $m$, and further the limit of $\{m\alpha\}$ for $m$ accepted is $\{-\beta\}$. Define $A_{4}$ to be the automaton that on input $n$ multiplies $n$ by the constant $d$ and then calls automaton $A_{3}$ on the result. Thus $A_{4}$ also accepts infinitely many $n$ and the limit of $\{dn\alpha\}=\{nb\gamma+cn\}=\{nb\gamma\}$ for $n$ accepted is still $\{-\beta\}$. Now define $A_{5}$ so it accepts if its input $r$ satisfies $r\equiv 0\ \mbox{\rm(mod}\ b\mbox{\rm)}$ and $A_{4}$ accepts $\lfloor r/b\rfloor$. Thus $A_{5}$ accepts the Ostrowski representations of an infinite sequence of natural numbers $r_{i}$ such that $\{r_{i}\gamma\}$ tends to $\{-\beta\}$. Hence, by the pumping lemma for regular languages, there exist finite words $x^{\prime},y^{\prime},z^{\prime}$, with $y^{\prime}$ nonempty, such that $A_{5}$ accepts $x^{\prime}{y^{\prime}}^{i}z^{\prime}$ for all $i\geq 0$. Define $x={z^{\prime}}^{R}$, $y={y^{\prime}}^{R}$, $z={x^{\prime}}^{R}$, where the exponent of $R$ denotes reversal.
Now we need an explicit expression for $\{m\gamma\}$ in terms of the continued fraction $a_{0},a_{1},\ldots]$ for an arbitrary real number $\gamma$, using two basic results on continued fractions that can be found, for example, in [44, Chap. 1]. First, if $p_{i}/q_{i}$ is the $i$’th convergent, we have $q_{i}\gamma-p_{i}=(-1)^{i}\gamma_{0}\gamma_{1}\cdots\gamma_{i}$, where $\gamma_{i}=[0,a_{i},a_{i+1},\ldots]$. Second, if the $\gamma$-Ostrowski representation of $m$ is $\sum_{0\leq i\leq t}e_{i}q_{i}$, then
$\{m\gamma\}=\sum_{0\leq i\leq t}e_{i}(q_{i}\gamma-p_{i})=\sum_{0\leq i\leq t}(-1)^{i}e_{i}\gamma_{0}\gamma_{1}\cdots\gamma_{i}.$
(Here we are using the “circular” representation of fractional parts where everything is understood to be taken mod $1$).
Now the particular $\gamma$ that concerns us has a purely periodic continued fraction expansion of period length $P$, so $\gamma_{i}=\gamma_{i+P}$ for all $i\geq 0$. So there is a constant $C<1$ such that $\gamma_{i}<C$ for all $i$.
Define $m_{j}$ to be the integer with least-significant-digit first Ostrowski representation given by $xy^{j}z$. Then from above we have
$\{m_{j}\gamma\}=\sum_{0\leq i\leq t}(-1)^{i}e_{i}\gamma_{0}\gamma_{1}\cdots\gamma_{i}$ (9)
where $e_{0}e_{1}\cdots e_{t}=xy^{j}z$, and further $\lim_{j\rightarrow\infty}\{m_{j}\gamma\}=\{-\beta\}$.
Let the infinite word $f_{0}f_{1}\cdots=xy^{\omega}=xyyy\cdots$. Consider the sum
$S=\sum_{i\geq 0}(-1)^{i}f_{i}\gamma_{0}\gamma_{1}\cdots\gamma_{i}.$ (10)
By comparison with Eq. (9) we see that $|S-\{m_{j}\gamma\}|=O(C^{|xy^{j}|})$, which is exponentially small in $j$. It follows that $S=\lim_{j\rightarrow\infty}\{m_{j}\gamma\}=\{-\beta\}$.
Now the terms $f_{i}$ are ultimately periodic with period $|y|$, and the $\gamma_{i}$ are periodic with period $P$. So Eq. (10) expresses $S$ as the sum of finitely many geometric series of the form $\sum_{i\geq 0}c_{\ell}\gamma_{\ell}^{P|y|i}$. Here each $c_{\ell}$ is a finite product of the $\gamma_{t}$, and hence in $\mathbb{Q}(\gamma)$. By the elementary formula for the sum of a geometric series, each series sum is in $\mathbb{Q}(\gamma)$, and hence so is $S=\{-\beta\}$ and hence so is $\beta$ itself. It is now easy to see that the specific expression of $\beta$ in terms of $\gamma$ is computable from the automaton. ∎